Twice the difference of seven times a number and four yields the same result as the same number increased by six.
*I am having a problem translating this problem, could you please help me*
Thank you!

Answer by checkley71(8403)

2(7x-4)=x+6
14x-8=x+6
14x-x=6+8
13x=14
x=14/13 answer.
proof:
2(7*14/13-4)=14/13+6
2(98/13)-8=(14+6*13)/13
196/13-8=(14+78)/13
(196-104)/13=92/13
92/13=92/13